OpenStudy (readergirl12):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle? ? in.

OpenStudy (dayakar):

let width =x in length = width+ 8 = x+8

OpenStudy (dayakar):

do u know perimeter formula

OpenStudy (readergirl12):

Um.. l+w?

OpenStudy (dayakar):

no 2(l+w)=p

OpenStudy (dayakar):

substitute above values in the equation can u do it

OpenStudy (dayakar):

2(x+8+x)=52

OpenStudy (readergirl12):

Ugh.. I was doing it in my head.. xD

OpenStudy (dayakar):

2(2x+8) =52

OpenStudy (readergirl12):

It is 12..?? o.e

OpenStudy (dayakar):

no show your work

OpenStudy (readergirl12):

Oh mer gerd.. This is so DIFFICULT .. >.<

OpenStudy (dayakar):

2x+8=52/2

OpenStudy (dayakar):

2x+8 =26 find x

OpenStudy (readergirl12):

Okie.. Subtract both sides from 8?